Opened 4 years ago

Last modified 3 years ago

#562 new flight processing

MA14/08, flight day 318/2014, North Selangor Peat Swamp Forest

Reported by: dap Owned by:
Priority: whenever Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~arsf/arsf_data/2014/flight_data/malaysia/MA14_08-2014_318_Selangor

Data arrived from ARSF via network transfer on 05/01/2015.

Scientific objective: to characterise spatially how oil palm cultivation is degrading tropical peatland ecosystem services, using airborne hyperspectral and LiDAR imagery, and supported by an intensive field campaign, in Malaysia’s North Selangor tropical peat swamp forest.

Priority:

PI: Paul Aplin

PI is happy to receive data by USB Hard Disk, network transfer or DVD
Co-I Doreen Boyd requested delivery via USB Hard Drive

OWL data were collected and are present.

Sensors:

  • Fenix (requested)
  • LiDAR (requested)
  • FW LiDAR (requested)

RCD images were requested but could not be collected due to an instrument failure.

Change History (25)

comment:1 Changed 4 years ago by stgo

Navigation processing

Had to split the basestation file using teqc. Command:
teqc -st 141113232000 -e 141114030000 +nav /tmp/318_output.14n Default_9488_0826_120823.m00 > /tmp/318_output.14o

basestation verification:

Lat 3 07 31.68451
Long 101 33 43.26145
height 18.941

comment:2 Changed 4 years ago by stgo

Hyperspectral processing
Found SCTs, waiting on calibration file to finalise processing.

Flightline FENIX
1 0.96
2 0.95
3 0.93
4 1.04
5 0.93
6 0.95
7 0.79

comment:3 Changed 4 years ago by stgo

Lidar Processing

Consistent pitch and roll found with values:

Pitch -0.00230
Roll -0.006225
Heading -0.00008

Classification complete, currently creating delivery.

comment:4 Changed 4 years ago by stgo

Lidar processing

Delivery creation complete, ready for DC

comment:5 Changed 4 years ago by gej

Started Lidar Delivery Check

comment:6 Changed 4 years ago by gej

Lidar Delivery Check

AGC values for FW data not saved (user_data field are 0 for min/max)

Everything else okay.

comment:7 Changed 4 years ago by stgo

Lidar Delivery Fix

I've replaced the discrete and full waveform files with updated ones including the AGC value, this should be everything that's needed I think.

Ready for redelivery check

comment:8 Changed 4 years ago by gej

Starting ReDelivery Check

comment:9 Changed 4 years ago by gej

Lidar Delivery Check

Pitch and Roll values appear incorrect. Were fine before updating to include AGC values.
Not sure what has caused this, as same Pitch and Roll values were used when reprocessing with the AGC saved.

comment:10 Changed 4 years ago by stgo

Lidar Processing
Reprocessed with new pitch roll values of:

Pitch -0.00260
Roll -0.00626

Files now include AGC value in point ID and user data field.

comment:11 Changed 4 years ago by gej

Lidar Delivery check

Read Me
Need to specify N/S in projection.

Renaming
Need to rename LAS1.0 folder to LAS1.2
Rename sol file in navigation

comment:12 Changed 4 years ago by stgo

Lidar Delivery

Fixed projection in readme

Renamed sol and las folder

comment:13 Changed 4 years ago by gej

Lidar Delivery Check

All looks good now,

Ready for Delivery

comment:14 Changed 4 years ago by gej

Lidar Delivered to Paul Aplin via FTP on 30th september 2015

comment:15 Changed 4 years ago by dac

  • Description modified (diff)

comment:16 Changed 4 years ago by gej

Lidar Delivered to Dr Doreen S. Boyd via USB hard disk on 5th October 2015 as requested.

comment:17 Changed 4 years ago by stgo

Fenix reprocessing

Ready for delivery check.

This delivery needs to be thoroughly checked for:

  • correct lines included
  • correct linenames from logsheet
  • data quality comments
  • spectra shifts

comment:18 Changed 4 years ago by gej

Fenix Delivery check

Screenshots need renaming,
Lines 5,6,7 have quite high cloud cover. should be in Readme

Data quality report needs updating to 2015 version
SCT's look okay,
Spectra looks okay.

Last edited 4 years ago by gej (previous) (diff)

comment:19 Changed 4 years ago by stgo

Fenix dc

fixed screenshot names

Data quality updated + name in project xml

Readme now has comment about cloud.

Last edited 4 years ago by stgo (previous) (diff)

comment:20 Changed 4 years ago by gej

Fenix DC
Screenshot names now fine,
Data quality report and name in projxml are both fine now.
Readme comment looks good.

Spectra needs checking by either lah or dac.

comment:21 Changed 4 years ago by lah

Fenix DC - spectra
Checked spectra of all lines and no major problems. There is a persistent spike in the overlap region between the VNIR and SWIR bands, which is fairly small and narrow, though occasionally it becomes quite an intense spike particularly when the overall signal is low. The spectral shift is within 5 nm compared to previous data (2014 169).

Had a quick check of the rest of the delivery and all looks fine, so zipping files. (Not so keen on pick fields in the screenshots though)

comment:22 Changed 4 years ago by gej

Fenix Delivered via FTP to Paul Aplin and via Hard drive to Doreen Boyd(University of Nottingham) on 11/11/2015

comment:23 Changed 4 years ago by dac

Sent hard drive to Paul Aplin on 23/11/2015 due to problems downloading data via FTP (slow transfer speed).

comment:24 Changed 3 years ago by dac

Archiving

Checked and ready for upload to NEODC.

comment:25 Changed 3 years ago by dac

Archiving

Uploaded to NEODC.

Note: See TracTickets for help on using tickets.